Scores of 59, 84, 61, 81 on all tests, use an inequality to find the minimum test score on final to pass with an average of 69 or higher, the final test counts as two tests
Answer by dabanfield(803)

Scores of 59, 84, 61, 81 on all tests, use an inequality to find the minimum test score on final to pass with an average of 69 or higher, the final test counts as two tests
Let x be the last test score. Since the average of the scores needs to be greater than or equal to 69 we know:
(59+84+61+81+x+x)/6 >= 69
Multiply both sides of the inequality by 6:
59+84+61+81+x+x >= 6*69
59+84+61+81+2*x >= 414
Substract 59+84+61+81 from both sides:
2*x >= 345-59-84-61-81
x >= ? (finish the calculations)